Entries for Selected Corporate Transactions Morrow Enterprises Inc. manufactures bathroom fixtures. Morrow Enterprises stockApr. 10. Issued 115,000 shares of common stock for $2,070,000. Debit Credit Date Account Apr. 10 Cash 2,070,000 Common StockDec. 28. Declared a $0.18-per-share dividend on common stock. Account Debit Credit Date Cash Dividends 712,750.0 X Dec. 28 Ca4. Prepare the Stockholders Equity section of the December 31, 20Y6, balance sheet. Morrow Enterprises Inc. Balance Sheet De
Answer #1

1)

Common stock
Jan. 1 Bal. $6000000
Apr 10 1150000
Aug 15 357500
Dec 31 Bal. $7507500
Paid-in-capital in excess of Stated value- Common stock
Jan 1 Bal. $1150000
Apr 10 920000
July 5 357500
Dec 31 Bal. $2427500
Retained Earnings
Dec 31 843295 Jan 1 Bal. $13620000
Dec 31 14165000
Dec 31 Bal. $26941705
Treasury Stock
Jan 1 Bal. $840000 June 6 840000
Nov 23 722000
Dec 31 Bal. $722000
Paid in Capital from sale of Treasury stock
June 6 180000
Stock Dividends Distributable
Aug 15 357500 July 5 357500
Stock Dividends
July 5 715000 Dec 31 715000
Cash Dividends
Dec 28 128295 Dec 31 128295

2)

July 5

Date Account Debit Credit
July 5 Stock Dividends (715000*5%*$20) $715000
Stock Dividends Distributable (715000*5%*$10) $357500
Paid in capital in excess of stated value- Common stock (715000-357500) $357500
(To record stock dividend declared)

Shares outstanding for stock dividend= 600000-60000+115000+60000= 715000 shares

Dec 28

Date Account Debit Credit
Dec 28 Cash Dividends (712750*$0.18) $128295
Cash Dividends Payable $128295
(To record dividend declared on common stock)

Shares outstanding for cash dividend= 600000-60000+115000+60000+35750-38000= 712750 shares

Dec 31

Date Account Debit Credit
Dec 31 Retained earnings (715000+128295) $843295
Stock dividends $715000
Cash dividends $128295
(To close stock dividend and cash dividend to retained earnings)

3)

Morrow Enterprises Inc.
Retained Earnings Statement
For the Year Ended December 31, 20Y6
Retained earnings, January 1, 20Y6 $13620000
Net income $14165000
Cash dividends (128295)
Stock dividends (715000)
Change in retained earnings 13321705
Retained earnings, December 31, 20Y6 $26941705

4)

Morrow Enterprises Inc.
Balance Shhet
December 31, 20Y6
Stockholders' equity
Paid-in-capital:
Common stock- $10 stated value (900000 shares authorized, 750750 shares issued) $7507500
Excess of issue price over stated value 2427500
From sale of treasury stock 180000
Total paid-in-capital 10115000
Retained earnings 26941705
Total 37056705
Treasury stock (38000 shares, at cost) (722000)
Total Stockholders' Equity $36334705